# Pre-processing and Compression: Understanding Hidden Representation Refinement Across Imaging Domains via Intrinsic Dimension

### v2

@@ Introduction
In this work, we first explore how these layer-by-layer representation ID evolution curves change for datasets beyond ImageNet, both for models trained on three other natural image datasets and for those trained on seven medical image datasets for a wide range of diagnostic tasks. We find that these curves have typically different shapes depending on if the model was trained on natural or medical images. In particular, the depth at which the ID peaks at some $d_{\mathrm{repr}}^{{\mathrm{max}}}$ is noticeably deeper into the model on average for natural image models, which we attribute to the task-relevant features of medical images typically requiring less abstract representations to capture, and so can be “pre-processed” for compression earlier in the network. Next, we explore the relationship between $d_{\mathrm{repr}}^{{\mathrm{max}}}$ and the intrinsic dimension of the raw data in its native space, $d_{\mathrm{data}}$ [19], and discover a new result: there is a strong correlation between the two across all datasets and models, showing that the maximum intrinsic complexity of learned representations in a model is guided by the intrinsic complexity of the model’s training data.
@@ 2 Experiments and Results
Evolution of hidden representation intrinsic dimension through network depth.
@@
Figure 1: Evolution of representation intrinsic dimension $d_{\mathrm{repr}}^{i}$ with respect to relative network depth $i/L$, for all models trained on each dataset.
@@ A.1 Additional Results
A.1.1 Evolution of hidden representation intrinsic dimension through network depth.
@@
Figure 3: Evolution of hidden representation intrinsic dimension $d_{\mathrm{repr}}^{i}$ with respect to relative network depth $i/L$, for all models trained on each dataset, for $N=500$. Top row: natural image models. Lower rows: medical image models.
Figure 4: Evolution of hidden representation intrinsic dimension $d_{\mathrm{repr}}^{i}$ with respect to relative network depth $i/L$, for all models trained on each dataset, for $N=750$. Top row: natural image models. Lower rows: medical image models.
Figure 5: Evolution of hidden representation intrinsic dimension $d_{\mathrm{repr}}^{i}$ with respect to relative network depth $i/L$, for all models trained on each dataset, for $N=1000$. Top row: natural image models. Lower rows: medical image models.
Figure 6: Evolution of hidden representation intrinsic dimension $d_{\mathrm{repr}}^{i}$ with respect to relative network depth $i/L$, for all models trained on each dataset, for $N=1250$. Top row: natural image models. Lower rows: medical image models.
Figure 7: Evolution of hidden representation intrinsic dimension $d_{\mathrm{repr}}^{i}$ with respect to relative network depth $i/L$, for all models trained on each dataset, for $N=1500$. Top row: natural image models. Lower rows: medical image models.

### v3

@@ Introduction
In this work, we first explore how layer-by-layer representation ID progression differs for models trained on datasets beyond ImageNet, including both three other natural image datasets and seven medical image datasets for a wide range of diagnostic tasks. We find that these curves have typically different shapes depending on if the model was trained on natural or medical images. In particular, the depth at which the ID peaks at some $d_{\mathrm{repr}}^{{\mathrm{max}}}$ is noticeably deeper into the model on average for natural image models, which we attribute to the task-relevant features of medical images typically requiring less abstract representations to capture, and so can be “pre-processed” for compression earlier in the network. Next, we explore the relationship between $d_{\mathrm{repr}}^{{\mathrm{max}}}$ and the intrinsic dimension of the raw data in its native space, $d_{\mathrm{data}}$ [19], and discover a new result: there is a strong correlation between the two across all datasets and models, showing that the maximum intrinsic complexity of learned representations in a model is guided by the intrinsic complexity of the model’s training data.
@@ 2 Experiments and Results
Progression of hidden representation intrinsic dimension through network depth.
@@
Figure 1: Progression of representation intrinsic dimension $d_{\mathrm{repr}}^{i}$ with respect to relative network depth $i/L$, for all models trained on each dataset.
@@ A.1 Additional Results
A.1.1 Progression of hidden representation intrinsic dimension through network depth.
@@
Figure 3: Progression of hidden representation intrinsic dimension $d_{\mathrm{repr}}^{i}$ with respect to relative network depth $i/L$, for all models trained on each dataset, for $N=500$. Top row: natural image models. Lower rows: medical image models.
Figure 4: Progression of hidden representation intrinsic dimension $d_{\mathrm{repr}}^{i}$ with respect to relative network depth $i/L$, for all models trained on each dataset, for $N=750$. Top row: natural image models. Lower rows: medical image models.
Figure 5: Progression of hidden representation intrinsic dimension $d_{\mathrm{repr}}^{i}$ with respect to relative network depth $i/L$, for all models trained on each dataset, for $N=1000$. Top row: natural image models. Lower rows: medical image models.
Figure 6: Progression of hidden representation intrinsic dimension $d_{\mathrm{repr}}^{i}$ with respect to relative network depth $i/L$, for all models trained on each dataset, for $N=1250$. Top row: natural image models. Lower rows: medical image models.
Figure 7: Progression of hidden representation intrinsic dimension $d_{\mathrm{repr}}^{i}$ with respect to relative network depth $i/L$, for all models trained on each dataset, for $N=1500$. Top row: natural image models. Lower rows: medical image models.